OpenStudy (anonymous):

solve 8= i/ square root symbol x

OpenStudy (anonymous):

\[8=\sqrt{x}\]

OpenStudy (anonymous):

question also has the square root as part of a fraction 1/ but i dont know how to write it

OpenStudy (anonymous):

try playing around with the equation button some more and just go something like \[8=i \sqrt{x/1}\] or is that it?

OpenStudy (anonymous):

no not that

OpenStudy (anonymous):

\[8=1/\sqrt{x}\]

OpenStudy (anonymous):

thats the best i can come up with

myininaya (myininaya):

\[8=\frac{1}{\sqrt{x}} =>(8)^2=(\frac{1}{\sqrt{x}})^2=>64=\frac{1}{x}=> 64x=1 => x=\frac{1}{64}\]

myininaya (myininaya):

so you didn't have i in the question?